# HiCMAE: Hierarchical Contrastive Masked Autoencoder for Self-Supervised Audio-Visual Emotion Recognition

### v1

@@ Abstract
Audio-Visual Emotion Recognition (AVER) has garnered increasing attention in recent years for its critical role in creating emotion-ware intelligent machines. Previous efforts in this area are dominated by the supervised learning paradigm. Despite significant progress, supervised learning is meeting its bottleneck due to the longstanding data scarcity issue in AVER. Motivated by recent advances in self-supervised learning, we propose Hierarchical Contrastive Masked Autoencoder (HiCMAE), a novel self-supervised framework that leverages large-scale self-supervised pre-training on vast unlabeled audio-visual data to promote the advancement of AVER. Following prior arts in self-supervised audio-visual representation learning, HiCMAE adopts two primary forms of self-supervision for pre-training, namely masked data modeling and contrastive learning. Unlike them which focus exclusively on top-layer representations while neglecting explicit guidance of intermediate layers, HiCMAE develops a three-pronged strategy to foster hierarchical audio-visual feature learning and improve the overall quality of learned representations. Firstly, it incorporates hierarchical skip connections between the encoder and decoder to encourage intermediate layers to learn more meaningful representations and bolster masked audio-visual reconstruction. Secondly, hierarchical cross-modal contrastive learning is also exerted on intermediate representations to narrow the audio-visual modality gap progressively and facilitate subsequent cross-modal fusion. Finally, during downstream fine-tuning, HiCMAE employs hierarchical feature fusion to comprehensively integrate multi-level features from different layers. To verify the effectiveness of HiCMAE, we conduct extensive experiments on 9 datasets covering both categorical and dimensional AVER tasks. Experimental results show that our method significantly outperforms state-of-the-art supervised and self-supervised audio-visual methods, which indicates that HiCMAE is a powerful audio-visual emotion representation learner. Codes and models will be publicly available at https://github.com/sunlicai/HiCMAE.
@@ 1 Introduction
Emotions are fundamental to the multifaceted spectrum of human experience, influencing our cognition, decision making, and interpersonal interactions schwarz2000emotion . They also play a pivotal role in developing intelligent machines and achieving the ultimate goal of emotional artificial intelligence, as Marvin Minsky, a pioneer of artificial intelligence, highlighted above minsky1988society . Typically, people express and perceive emotions through multiple modalities. Previous psychological studies have demonstrated that the language modality (i.e., verbal information) only contributes 7% to the perception of emotions in our daily communication, while the audio (e.g., tone and intonation) and visual (e.g., facial expressions) modalities significantly predominate, contributing to 38% and 55% respectively mehrabian1968communication ; wu2014survey . As a result, the last two decades have witnessed a surge of interest in automatic Audio-Visual Emotion Recognition (AVER) in the affective computing community.
@@
Early studies on AVER concentrate on developing various hand-engineering features for audio and video modalities zeng2008survey ; wu2014survey . With the advent of the deep learning era, a new trend has emerged towards learning features directly from raw audio-visual data by training deep supervised neural networks in an end-to-end manner zhang2017learning ; tzirakis2017end ; zhang2023transformer . Despite considerable advancements, supervised learning is heavily constrained by its reliance on large amounts of labeled data to achieve satisfactory performance. This reliance significantly hampers further progress of supervised methods due to the longstanding issue of data scarcity in AVER.
@@
Recently, another deep learning paradigm, i.e., self-supervised learning, which can learn powerful representations from vast unlabeled data, has revolutionized many research areas balestriero2023cookbook . In particular, masked data modeling (e.g., MAE he2022masked ) and contrastive learning (e.g., CLIP radford2021learning ) are two prominent methods and have demonstrated great success in self-supervised visual representation learning. Specifically, masked data modeling aims to reconstruct the raw data from masked inputs, while contrastive learning encourages the semantic alignment of different modalities. They are also combined and extended to learn generic audio-visual representations in recent studies, such as CAV-MAE gong2023contrastive and MAViL huang2023mavil . However, due to the domain gap between upstream pre-training and downstream AVER tasks, the learned representations are typically not suitable for AVER. More recently, Sadok et al. sadok2023vector developed a vector-quantized masked autoencoder (VQ-MAE-AV) for AVER. Although achieving promising results, VQ-MAE-AV requires pre-trained variational autoencoders for vector quantization and thus cannot be trained in a single stage. More importantly, several studies on masked image modeling have shown that intermediate layers are essential to self-supervised visual representation learning and explicitly guiding them can improve the quality of learned representations wang2023masked ; liu2023improving . However, the aforementioned methods fail to achieve that goal as they solely operate on representations from the top layer and neglect explicit guidance on intermediate layers, thus impeding feature evolution across layers and leading to sub-optimal results in downstream tasks.
@@
Figure 1: Comparison with state-of-the-art audio-visual methods on 9 datasets. We present Pearson correlation coefficient on Werewolf-XL zhang2021werewolf and weighted F1-score on AVCAffe sarkar2022avcaffe and MER-MULTI lian2023mer . For other datasets, we show weighted average recall (WAR).
@@
To address the above challenges, this paper builds on top of MAE and contrastive learning to propose Hierarchical Contrastive Masked Autoencoder (HiCMAE), a novel self-supervised framework tailored for AVER. As shown in Fig. 2, it is mainly composed of two audio-visual encoders, a cross-modal fusion encoder, and two lightweight audio-visual decoders. To foster hierarchical audio-visual feature learning and improve the overall quality of learned representations, HiCMAE develops a three-pronged strategy: (1) Drawing inspiration from the architecture design in U-Net ronneberger2015u , HiCMAE introduces hierarchical skip connections between the encoder and decoder to drive intermediate encoder layers to learn more useful representations and aid the decoder in accomplishing the task of masked audio-visual reconstruction. (2) Considering the natural audio-visual correspondences in videos, hierarchical cross-modal contrastive learning is also applied to intermediate representations of audio-visual encoders to reduce heterogeneous modality gap in a progressive manner and enhance cross-modal fusion in subsequent layers. (3) Since different layers typically capture distinct levels of information, HiCMAE performs hierarchical feature fusion during downstream fine-tuning to comprehensively integrate multi-level features from various encoder layers. To demonstrate the effectiveness of HiCMAE, we perform large-scale self-supervised pre-training on VoxCeleb2 chung2018voxceleb2 and evaluate the pre-trained model on nine datasets encompassing both categorical and dimensional AVER tasks. As illustrated in Fig. 1, our HiCMAE significantly outperforms state-of-the-art supervised or self-supervised audio-visual methods. For example, HiCMAE surpasses the best-performing VQ-MAE-AV sadok2023vector by +4.49% WAR on CREMA-D (6-class) cao2014crema and beats the state-of-the-art T-MEP zhang2023transformer by +5.02% WAR on MAFW (11-class) liu2022mafw and +6.16% WAR on DFEW jiang2020dfew . Moreover, extensive ablation studies also justify various design choices in HiCMAE.
@@
    Unlike previous methods, HiCMAE introduces a three-pronged approach (i.e., hierarchical skip connections, hierarchical cross-modal contrastive learning, and hierarchical feature fusion) to foster hierarchical audio-visual feature learning and the ablation studies verify its efficacy.
@@ 2.1 Audio-Visual Emotion Recognition
Most studies on Audio-Visual Emotion Recognition (AVER) fall into the supervised learning paradigm. They mainly focus on two important aspects: unimodal feature extraction and audio-visual information fusion. As for the first aspect, researchers have developed and exploited numerous features in the past two decades zeng2008survey ; wu2014survey ; zhang2023deep ; lian2024merbench . Early studies concentrate on various handcrafted features for two modalities, such as IS13 schuller2013interspeech and eGeMAPS eyben2015geneva for audio, LBP-TOP zhao2007dynamic and HOG dalal2005histograms for video. With the advent of deep learning, a large amount of deep supervised models trained on large audio and image/video datasets have emerged as powerful audio-visual feature extractors fan2016video ; chen2017multimodal ; sun2020multi ; sun2021multimodal ; meng2022valence ; lian2023mer , such as PANNs kong2020panns and VGGish hershey2017cnn for audio, VGGFace cao2018vggface2 and C3D tran2015learning for video. There are also lots of attempts to train end-to-end deep neural networks on raw audio and video emotion data trigeorgis2016adieu ; tzirakis2017end ; huang2018end ; jiang2020dfew ; wang2022ferv39k . In recent years, plenty of large self-supervised pre-trained models have demonstrated great success in audio (e.g., Wav2vec2.0 baevski2020wav2vec and HuBERT hsu2021hubert ) or video (e.g., SVFAP sun2023svfap and MAE-DFER sun2023mae ) emotion recognition. After unimodal feature extraction, the next crucial step is audio-visual information fusion. Current fusion strategies can be roughly divided into three categories, i.e., early fusion, late fusion, and model-level fusion zeng2008survey ; wu2014survey . Early fusion typically combines audio-visual features at the input level chen2017multimodal ; meng2022valence . In contrast, late fusion integrates audio-visual predictions at the decision level sun2020multi ; sun2021multimodal . The most commonly used strategy is model-level fusion zhang2017learning ; tsai2019multimodal ; huang2020multimodal ; goncalves2022auxformer ; lian2021ctnet ; tran2022pre ; sun2023efficient ; zhang2023transformer . For example, MulT tsai2019multimodal utilizes cross-modal attention to capture dense interactions in unaligned audio-visual feature sequences. EMT sun2023efficient improves MulT by introducing the global multimodal context and employing it to interact with local unimodal features to achieve efficient cross-modal information exchange. The recent T-MEP zhang2023transformer adopts a similar strategy with MulT to fuse fine-grained audio-visual tokens by combining self-attention and cross-attention mechanisms in an interleaving manner.
@@ 2.2 Self-Supervised Audio-Visual Representation Learning
In general, there are three kinds of methods for learning generic audio-visual representations through self-supervision, namely contrastive learning, masked data modeling, and the hybrid method. Contrastive learning exploits the natural audio-visual correspondence in videos as a free signal for self-supervision arandjelovic2017look ; owens2018audio ; morgado2021audio . The goal of masked data modeling is to reconstruct the original data from its masked input. Motivated by its recent success in the image and video domain he2022masked ; tong2022videomae ; feichtenhofer2022masked , Georgescu et al. georgescu2023audiovisual extended it to the audio-visual domain and achieve significant improvements over previous methods. Recently, a few studies have made attempts to combine the former two kinds of methods, resulting in the hybrid method. In particular, CAV-MAE gong2023contrastive integrates MAE and cross-modal contrastive learning and shows that they are complementary. MAViL huang2023mavil further introduces intra-modal contrastive learning and masked self-training on contextualized features to improve audio-visual pre-training. Although these methods have demonstrated great success in general audio-visual tasks, the learned representations are typically not suitable for AVER because they are trained on general scene or action videos instead of facial videos in AVER. Recently, VQ-MAE-AV sadok2023vector introduces a vector-quantized MAE for AVER. Despite promising results, it requires two-stage training. In addition, the more important issue is that the aforementioned audio-visual masked autoencoders fail to promote hierarchical audio-visual feature learning as they focus exclusively on top-layer representations while neglecting explicit guidance of intermediate layers, thus impeding feature evolution across layers and leading to sub-optimal performance wang2023masked ; liu2023improving . Therefore, our HiCAME presents a three-pronged approach to foster hierarchical audio-visual feature learning and demonstrates improved performance.
@@ 3 Method
In this section, we elaborate on Hierarchical Contrastive Masked AutoEncoder (HiCMAE), a novel self-supervised audio-visual representation learning framework for AVER. The whole pre-training pipeline of HiCMAE is illustrated in Fig. 2. It mainly consists of two modality-specific encoders, a cross-modal fusion encoder, and two lightweight modality-specific decoders. HiCMAE adopts two primary forms of self-supervision for pre-training: masked data modeling (i.e., masked audio-visual reconstruction) and contrastive learning. Moreover, it introduces a three-pronged strategy to promote hierarchical audio-visual feature learning during both pre-training and fine-tuning, including hierarchical skip connections between the encoder and decoder, hierarchical cross-modal contrastive learning, and hierarchical feature fusion for downstream fine-tuning.
@@ 3.1 Masked Audio-Visual Reconstruction with Hierarchical Skip Connections
As depicted in Fig. 2, HiCMAE follows MAE he2022masked to adopt an asymmetric encoder-decoder architecture for efficient self-supervised audio-visual pre-training. The raw audio and video inputs are first embedded into tokens and then masked to filter out a substantial proportion of tokens. Next, the visible (i.e., unmasked) audio and video tokens are processed by their modality-specific encoders and a cross-modal fusion encoder. After feature encoding, the visible audio and video tokens are padded with learnable masked tokens and then passed through lightweight modality-specific decoders for final audio and video reconstruction. Note that, unlike conventional masked autoencoders, hierarchical skip connections are added between audio-visual encoders and decoders to better guide the encoder feature learning at different levels and promote masked audio-visual reconstruction by providing the decoder with multi-level features.
@@ 3.1.1 Data Embedding and Token Masking
Since we use Transformer vaswani2017attention as the main component of the encoder, it is necessary to embed audio and video inputs into a sequence of discrete tokens. Formally, we denote the facial video frames as $\mathbf{X}_{v}\in\mathbb{R}^{T_{v}\times H\times W\times 3}$ ($T_{v}$ is the number of frames, $H$ and $W$ denote the height and width, and $3$ represents RGB channels) and the audio spectrogram as $\mathbf{X}_{a}\in\mathbb{R}^{T_{a}\times F}$ ($T_{a}$ is the temporal length and $F$ denotes the number of frequency channels). We utilize a cube embedding layer with a size of $2\times 16\times 16$ to split $\mathbf{X}_{v}$ into video tokens $\mathbf{X}_{v}^{{}^{\prime}}\in\mathbb{R}^{N_{v}\times C}$ ($N_{v}=\frac{T}{2}\cdot\frac{H}{16}\cdot\frac{W}{16}$ is the number of video tokens, $C$ is the number of feature channels). As for $\mathbf{X}_{a}$, we employ a patch embedding layer with a size of $16\times 16$ to split it into audio tokens $\mathbf{X}_{a}^{{}^{\prime}}\in\mathbb{R}^{N_{a}\times C}$ ($N_{a}=\frac{T_{a}}{16}\cdot\frac{F}{16}$ is the number of audio tokens).
After data embedding, we mask out a large proportion of audio and video tokens to make audio-visual reconstruction a non-trivial self-supervised task and significantly reduce the pre-training cost at the same time. Considering the high temporal redundancy and correlation in video data, we adopt the tube masking (i.e., each temporal slice has the same masking pattern) strategy tong2022videomae . For audio, we simply use random masking huang2022masked . The masking ratios for audio and video tokens are set to $\rho_{a}=80\%$ and $\rho_{v}=90\%$ respectively tong2022videomae ; huang2022masked . After token masking, only the visible video tokens $\mathbf{X}_{v}^{{}^{\prime\prime}}\in\mathbb{R}^{N_{v}^{\prime}\times C}$ ($N_{v}^{\prime}=(1-\rho_{v})\cdot N_{v}$) and audio tokens $\mathbf{X}_{a}^{{}^{\prime\prime}}\in\mathbb{R}^{N_{a}^{\prime}\times C}$ ($N_{a}^{\prime}=(1-\rho_{a})\cdot N_{a}$) will be processed by the encoder introduced next.

### v2

@@ Abstract
Audio-Visual Emotion Recognition (AVER) has garnered increasing attention in recent years for its critical role in creating emotion-aware intelligent machines. Previous efforts in this area are dominated by the supervised learning paradigm. Despite significant progress, supervised learning is meeting its bottleneck due to the longstanding data scarcity issue in AVER. Motivated by recent advances in self-supervised learning, we propose Hierarchical Contrastive Masked Autoencoder (HiCMAE), a novel self-supervised framework that leverages large-scale self-supervised pre-training on vast unlabeled audio-visual data to promote the advancement of AVER. Following prior arts in self-supervised audio-visual representation learning, HiCMAE adopts two primary forms of self-supervision for pre-training, namely masked data modeling and contrastive learning. Unlike them which focus exclusively on top-layer representations while neglecting explicit guidance of intermediate layers, HiCMAE develops a three-pronged strategy to foster hierarchical audio-visual feature learning and improve the overall quality of learned representations. Firstly, it incorporates hierarchical skip connections between the encoder and decoder to encourage intermediate layers to learn more meaningful representations and bolster masked audio-visual reconstruction. Secondly, hierarchical cross-modal contrastive learning is also exerted on intermediate representations to narrow the audio-visual modality gap progressively and facilitate subsequent cross-modal fusion. Finally, during downstream fine-tuning, HiCMAE employs hierarchical feature fusion to comprehensively integrate multi-level features from different layers. To verify the effectiveness of HiCMAE, we conduct extensive experiments on 9 datasets covering both categorical and dimensional AVER tasks. Experimental results show that our method significantly outperforms state-of-the-art supervised and self-supervised audio-visual methods, which indicates that HiCMAE is a powerful audio-visual emotion representation learner. Codes and models are publicly available at https://github.com/sunlicai/HiCMAE.
@@ 1 Introduction
Emotions are fundamental to the multifaceted spectrum of human experience, influencing our cognition, decision-making, and interpersonal interactions [1]. They also play a pivotal role in developing intelligent machines and achieving the ultimate goal of emotional artificial intelligence, as Marvin Minsky, a pioneer of artificial intelligence, highlighted above [2]. Typically, people express and perceive emotions through multiple modalities. Previous psychological studies have demonstrated that the language modality (i.e., verbal information) only contributes 7% to the perception of emotions in our daily communication, while the audio (e.g., tone and intonation) and visual (e.g., facial expressions) modalities significantly predominate, contributing to 38% and 55% respectively [3, 4]. As a result, the last two decades have witnessed a surge of interest in automatic Audio-Visual Emotion Recognition (AVER) in the affective computing community. The task of AVER is to extract emotion-related representations from audio-visual signals and then integrate them to identify the subject’s emotional state. Several concrete examples are depicted in Fig. 1.
Figure 1: Several samples selected from the MAFW [5] dataset.
@@
Early studies on AVER concentrate on developing various hand-engineering features for audio and video modalities [6, 4]. With the advent of the deep learning era, a new trend has emerged towards learning features directly from raw audio-visual data by training deep supervised neural networks in an end-to-end manner [7, 8, 9]. Despite considerable advancements, supervised learning is heavily constrained by its reliance on large amounts of labeled data to achieve satisfactory performance. This reliance significantly hampers further progress of supervised methods due to the longstanding issue of data scarcity in AVER [10, 11].
@@
Recently, another deep learning paradigm, i.e., self-supervised learning, which can learn powerful representations from vast unlabeled data, has revolutionized many research areas [12, 13, 14, 15, 16, 17]. In particular, masked data modeling (e.g., MAE [17]) and contrastive learning (e.g., CLIP [16]) are two prominent methods and have demonstrated great success in self-supervised visual representation learning. Specifically, masked data modeling aims to reconstruct the raw data from masked inputs, while contrastive learning encourages the semantic alignment of different modalities. They are also combined and extended to learn generic audio-visual representations in recent studies, such as CAV-MAE [18] and MAViL [19]. However, due to the domain gap between upstream pre-training and downstream AVER tasks, the learned representations are typically not suitable for AVER. More recently, Sadok et al. [20] developed a vector-quantized masked autoencoder (VQ-MAE-AV) for AVER. Although achieving promising results, VQ-MAE-AV requires pre-trained variational autoencoders for vector quantization and thus cannot be pre-trained in a single stage. More importantly, several studies on masked image modeling have shown that intermediate layers are essential to self-supervised visual representation learning and explicitly guiding them can improve the quality of learned representations [21, 22]. However, the aforementioned methods fail to achieve that goal as they solely operate on representations from the top layer and neglect explicit guidance on intermediate layers, thus impeding feature evolution across layers and leading to sub-optimal results in downstream tasks.
@@
Figure 2: Comparison with state-of-the-art audio-visual methods on 9 datasets. We present Pearson correlation coefficient on Werewolf-XL [23] and weighted F1-score on AVCAffe [24] and MER-MULTI [25]. For other datasets, we show weighted average recall (WAR).
@@
To address the above challenges, this paper builds on top of MAE and contrastive learning to propose Hierarchical Contrastive Masked Autoencoder (HiCMAE), a novel self-supervised framework tailored for AVER. As shown in Fig. 3, it is mainly composed of two audio-visual encoders, a cross-modal fusion encoder, and two lightweight audio-visual decoders. To foster hierarchical audio-visual feature learning and improve the overall quality of learned representations, HiCMAE develops a three-pronged strategy: (1) Drawing inspiration from the architecture design in U-Net [26], HiCMAE introduces hierarchical skip connections between the encoder and decoder to drive intermediate encoder layers to learn more useful representations and aid the decoder in accomplishing the task of masked audio-visual reconstruction. (2) Considering the natural audio-visual correspondences in videos, hierarchical cross-modal contrastive learning is also applied to intermediate representations of audio-visual encoders to reduce heterogeneous modality gap in a progressive manner and enhance cross-modal fusion in subsequent layers. (3) Since different layers typically capture distinct levels of information, HiCMAE performs hierarchical feature fusion during downstream fine-tuning to comprehensively integrate multi-level features from various encoder layers. To demonstrate the effectiveness of HiCMAE, we perform large-scale self-supervised pre-training on VoxCeleb2 [27] and evaluate the pre-trained model on nine datasets encompassing both categorical and dimensional AVER tasks. As illustrated in Fig. 2, our HiCMAE significantly outperforms state-of-the-art supervised or self-supervised audio-visual methods. For example, HiCMAE surpasses the best-performing VQ-MAE-AV [20] by +4.49% WAR on CREMA-D (6-class) [28] and beats the state-of-the-art T-MEP [9] by +5.02% WAR on MAFW (11-class) [5] and +6.16% WAR on DFEW [29]. Moreover, extensive ablation studies also justify various design choices in HiCMAE.
@@
    Unlike previous methods, HiCMAE introduces a three-pronged approach to foster hierarchical audio-visual feature learning and the ablation studies verify its efficacy.
@@ 2.1 Audio-Visual Emotion Recognition
Most studies on Audio-Visual Emotion Recognition (AVER) fall into the supervised learning paradigm. They mainly focus on two important aspects: unimodal feature extraction and audio-visual information fusion. As for the first aspect, researchers have developed and exploited numerous features in the past two decades [6, 4, 10, 30]. Early studies concentrate on various handcrafted features for two modalities, such as IS13 [31] and eGeMAPS [32] for audio, LBP-TOP [33] and HOG [34] for video. With the advent of deep learning, a large amount of deep supervised models trained on large audio and image/video datasets have emerged as powerful audio-visual feature extractors [35, 36, 37, 38, 39, 25], such as PANNs [40] and VGGish [41] for audio, VGGFace [42] and C3D [43] for video. There are also lots of attempts to train end-to-end deep neural networks on raw audio and video emotion data [44, 8, 45, 29, 46]. In recent years, plenty of large self-supervised pre-trained models have demonstrated great success in audio (e.g., Wav2vec2.0 [47] and HuBERT [48]) or video (e.g., SVFAP [49] and MAE-DFER [50]) emotion recognition. After unimodal feature extraction, the next crucial step is audio-visual information fusion. Current fusion strategies can be roughly divided into three categories, i.e., early fusion, late fusion, and model-level fusion [6, 4]. Early fusion typically combines audio-visual features at the input level [36, 39]. In contrast, late fusion integrates audio-visual predictions at the decision level [37, 38]. The most commonly used strategy is model-level fusion [7, 51, 52, 53, 54, 55, 56, 57, 9]. For example, MulT [51] utilizes cross-modal attention to capture dense interactions in unaligned audio-visual feature sequences. EMT [57] improves MulT by introducing the global multimodal context and employing it to interact with local unimodal features to achieve efficient cross-modal information exchange. The recent T-MEP [9] adopts a similar strategy with MulT to fuse fine-grained audio-visual tokens by combining self-attention and cross-attention mechanisms in an interleaving manner.
@@ 2.2 Self-Supervised Audio-Visual Representation Learning
In general, there are three kinds of methods for learning generic audio-visual representations through self-supervision, namely contrastive learning, masked data modeling, and the hybrid method. Contrastive learning exploits the natural audio-visual correspondence in videos as a free signal for self-supervision [58, 59, 60]. In contrast, the goal of masked data modeling is to reconstruct the original data from its masked input. Motivated by its recent success in the image and video domain [17, 61, 62], Georgescu et al. [63] extend it to the audio-visual domain and achieve significant improvements over previous methods. Recently, a few studies have made attempts to combine the former two kinds of methods, resulting in the hybrid method. In particular, CAV-MAE [18] integrates MAE and cross-modal contrastive learning and shows that they are complementary. MAViL [19] further introduces intra-modal contrastive learning and masked self-training on contextualized features to improve audio-visual pre-training. Although these methods have demonstrated great success in general audio-visual tasks, the learned representations are typically not suitable for AVER because they are trained on general scene or action videos instead of facial videos in AVER. Recently, VQ-MAE-AV [20] introduces a vector-quantized MAE for AVER. Despite promising results, it requires two-stage pre-training. In addition, the more important issue is that the aforementioned audio-visual masked autoencoders fail to promote hierarchical audio-visual feature learning as they focus exclusively on top-layer representations while neglecting explicit guidance of intermediate layers, thus impeding feature evolution across layers and leading to sub-optimal performance [21, 22]. Therefore, our HiCAME presents a three-pronged approach to foster hierarchical audio-visual feature learning and demonstrates improved performance.
@@ 3 Method
In this section, we elaborate on Hierarchical Contrastive Masked AutoEncoder (HiCMAE), a novel self-supervised audio-visual emotion representation learning framework for AVER. The training process of HiCMAE includes two steps, i.e., self-supervised pre-training (Section 3.1-3.3) on large-scale unlabeled AVER data and downstream fine-tuning (Section 3.4) on limited labeled AVER data. Specifically, the self-supervised pre-training pipeline of HiCMAE is illustrated in Fig. 3. It mainly consists of two modality-specific encoders, a cross-modal fusion encoder, and two lightweight modality-specific decoders. HiCMAE adopts two primary forms of self-supervision for pre-training: masked data modeling (i.e., masked audio-visual reconstruction) and contrastive learning. Moreover, it introduces a three-pronged strategy to promote hierarchical audio-visual feature learning during both pre-training and fine-tuning, including hierarchical skip connections between the encoder and decoder (Section 3.1), hierarchical cross-modal contrastive learning (Section 3.2), and hierarchical feature fusion for downstream fine-tuning (Section 3.4).
@@ 3.1 Masked Audio-Visual Reconstruction with Hierarchical Skip Connections
As depicted in Fig. 3, HiCMAE follows MAE [17] to adopt an asymmetric encoder-decoder architecture for efficient self-supervised audio-visual pre-training. The raw audio and video inputs are first embedded into tokens and then masked to filter out a substantial proportion of tokens. Next, the visible (i.e., unmasked) audio and video tokens are processed by their modality-specific encoders and a cross-modal fusion encoder. After feature encoding, the visible audio and video tokens are padded with learnable masked tokens and then passed through lightweight modality-specific decoders for final audio and video reconstruction. Note that, unlike conventional masked autoencoders, hierarchical skip connections are added between audio-visual encoders and decoders to better guide the encoder feature learning at different levels and promote masked audio-visual reconstruction by providing the decoder with multi-level features.
@@ 3.1.1 Data Embedding and Token Masking
Since we use Transformer [64] as the main component of the encoder, it is necessary to embed audio and video inputs into a sequence of discrete tokens. Formally, we denote the facial video frames as $\mathbf{X}_{v}\in\mathbb{R}^{T_{v}\times H\times W\times 3}$ ($T_{v}$ is the number of frames, $H$ and $W$ denote the height and width, and $3$ represents RGB channels) and the audio spectrogram as $\mathbf{X}_{a}\in\mathbb{R}^{T_{a}\times F}$ ($T_{a}$ is the temporal length and $F$ denotes the number of frequency channels). We utilize a cube embedding layer with a size of $2\times 16\times 16$ to split $\mathbf{X}_{v}$ into video tokens $\mathbf{X}_{v}^{{}^{\prime}}\in\mathbb{R}^{N_{v}\times C}$ ($N_{v}=\frac{T}{2}\cdot\frac{H}{16}\cdot\frac{W}{16}$ is the number of video tokens, $C$ is the number of feature channels). As for $\mathbf{X}_{a}$, we employ a patch embedding layer with a size of $16\times 16$ to split it into audio tokens $\mathbf{X}_{a}^{{}^{\prime}}\in\mathbb{R}^{N_{a}\times C}$ ($N_{a}=\frac{T_{a}}{16}\cdot\frac{F}{16}$ is the number of audio tokens).
After data embedding, we mask out a large proportion of audio and video tokens to make audio-visual reconstruction a non-trivial self-supervised task and significantly reduce the pre-training cost at the same time. Considering the high temporal redundancy and correlation in video data, we adopt the tube masking (i.e., each temporal slice has the same masking pattern) strategy [61]. For audio, we simply use random masking [65]. The masking ratios for audio and video tokens are set to $\rho_{a}=80\%$ and $\rho_{v}=90\%$ respectively [61, 65]. After token masking, only the visible video tokens $\mathbf{X}_{v}^{{}^{\prime\prime}}\in\mathbb{R}^{N_{v}^{\prime}\times C}$ ($N_{v}^{\prime}=(1-\rho_{v})\cdot N_{v}$) and audio tokens $\mathbf{X}_{a}^{{}^{\prime\prime}}\in\mathbb{R}^{N_{a}^{\prime}\times C}$ ($N_{a}^{\prime}=(1-\rho_{a})\cdot N_{a}$) will be processed by the encoder introduced next.